OpenStudy (anonymous):

For each of the following, decide if the given vector field is a gradient of a function f. If so, give the function f nd enter it as your answer; if not, enter the word NONE for your answer and be sure that you are able to explain why this is the case. (b) 6xcos(y^2)î + (2ycos((x^2) + (y^2)))j (c) [6cos((x^2) + (2y^2)) - 12(x^2)sin((x^2) + (2y^2))] - [24xysin((x^2) + (2y^2))]

OpenStudy (anonymous):

@mathmale HELP again??

OpenStudy (anonymous):

@Disco619 this is Calc 3 I hope you can do it!

OpenStudy (luigi0210):

@TuringTest

OpenStudy (anonymous):

@Snowfire Can you help me with this question?

OpenStudy (disco619):

It's either @Snowfire or @Denu8 . Denu is offline. Snow might be able to do this... i hope.

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (turingtest):

if a vector field is conservative, it is the gradient of some potential function so we need to know if these fields are conservative

OpenStudy (disco619):

@ranga I THINK YOU CAN DO THIS!

OpenStudy (turingtest):

a vector field \[\vec F=P\hat i+Q\hat j\]is conservative if\[{\partial P\over \partial y}={\partial Q\over\partial x}\]

OpenStudy (turingtest):

take those partials and compare them if they are equal, the vector field is the gradient of some potential function

OpenStudy (turingtest):

as far as finding the function, that takes a bit of integration the process is outlined nicely here in example 2: http://tutorial.math.lamar.edu/Classes/CalcIII/ConservativeVectorField.aspx

OpenStudy (anonymous):

Keep going @TuryingTest I still need help lol... I can't seem to get the partial of either of them. and what do i do AFTER i know they are conservative?

OpenStudy (turingtest):

if you can't get the partial then there's not much point in trying to fund a possibly non-existent potential function, but you would take note of the fact that by definition\[\vec F=\nabla f=\frac{\partial f}{\partial x}\hat i+\frac{\partial f}{\partial y}\hat j\implies P=\frac{\partial f}{\partial x}, ~~Q=\frac{\partial f}{\partial y}\]

OpenStudy (turingtest):

for the first one, 6xcos(y^2)î + (2ycos((x^2) + (y^2)))j P= 6xcos(y^2) what the partial of that wrt y?

OpenStudy (anonymous):

6xsin(y^2)(2)...?

OpenStudy (turingtest):

dropped a y

OpenStudy (anonymous):

Oh 6x(sin(y^2))(2y) right?

OpenStudy (turingtest):

right, not what's partial wrt x of Q= (2ycos((x^2) + (y^2)))

OpenStudy (turingtest):

now*

OpenStudy (anonymous):

2ycos(x^2)(2x) + (y^2)x ?

OpenStudy (turingtest):

erm, not if I'm reading those parentheses right is it\[Q=2y\cos(x^2+y^2)\]?

OpenStudy (anonymous):

yes that is what Q is

OpenStudy (turingtest):

so then treating y as constant...\[\frac{\partial Q}{\partial x}=-2y\sin(x^2+y^2)\frac{\partial }{\partial x}(x^2+y^2)=-2xy\sin(x^2+y^2)\]

OpenStudy (anonymous):

oh ok i think i was still doing wrt y like we did for P. OK so now what and seriously thanks so much for working with me here

OpenStudy (turingtest):

oh typo, should be a 4 at the end\[\frac{\partial Q}{\partial x}=-2y\sin(x^2+y^2)\frac{\partial }{\partial x}(x^2+y^2)=-4xy\sin(x^2+y^2)\]

OpenStudy (turingtest):

you're welcome anyway, is \[{\partial P\over \partial y}={\partial Q\over\partial x}\]?

OpenStudy (anonymous):

umm no?

OpenStudy (turingtest):

so then \(\vec F\) is not conservative, which means it has no potential function \(f\), so we're done

OpenStudy (turingtest):

I'll let you check if\[{\partial P\over \partial y}={\partial Q\over\partial x}\]for the next oneon your own

OpenStudy (anonymous):

@mathmale I need help with part C

OpenStudy (anonymous):

@TuringTest This is the work I have so far for part C. @mathmale could ya help me out?

OpenStudy (anonymous):

@ranga Can you help me on this?

OpenStudy (anonymous):

I NEED THIS QUESTION ANSWERED BY 5:00 TONIGHT!